Morris Dancing is a ritual type of dancing that flourished in the Cotswold hills of England. Based on a long tradition of English folk dance, it is derived from ancient ceremonies celebrating the seasons, particularly the solstices.

The dancers leap and bound in ritual patterns, waking the earth with the jingling of bells, which are worn on the dancers’ legs. The clashing of sticks is symbolic of chasing away evil spirits so that the crops may grow and prosper.

Shakespeare knew the Morris; Henry VIII was entertained by Morris dancers at Christmas time. Dancers typically performed on the greens and streets of their own towns and neighboring villages, entertaining at seasonal events and fairs, adding color, music and ritual.

Most of its origins and history are speculative. It is suggested that it began on the moors and was originally called “Moorish” dancing. Although not competitive, a group of Morris dancers is called a “team” because they must work together to form the finished routine. They are taught by their “Foreman,” business arrangements are handled by the “Squire,” and finances are handled by the “Bagman.”

Middlesex Morris, the oldest mixed Morris team in the Boston area, celebrates its tenth anniversary this year. Popular for their infectious good humor and high spirits as well as their dancing, the dancers and musicians present a wonderful show that will delight audiences of all ages.

If you would like Middlesex Morris to dance for your event or function, or if you are interested in dancing with us, contact one of the people listed below. Middlesex Morris welcomes new dancers each year in October.

Squire: Foreperson:

Annette Sassi Sue Jick 23 Sunset Rd. 19 Smith St. Somerville, MA 02144 Groton, MA 01450 (617) 625–2924 (508) 448–2497 asassi@edc.org

Middlesex Morris’s homepage on the World Wide Web:

                                                                 http://www.mit.edu:8001/people/jcb/Middlesex.html

Printed on recycled paper. Please recycle.

Middlesex Morris

in tree: Brian Wilson, Paul Balliet; standing: Lyn Mullen, Dan Spurr, Sue Jick, Charlie Spurr; kneeling: Forrest Larson; seated: Annette Sassi, Dan Groher, Jerry Callen, Jeff Bigler, Bill Bumpus

Middlesex Morris

10^{th} Anniversary Season 1994